# The relevance of degenerate states in chiral polaritonics

### v1

@@ IV Summary and conclusion
Nevertheless, our here-presented results suggest a remedy for this fundamental theoretical issue of chiral polaritonics. Instead of using numerically challenging and mathematically problematic (not every choice is consistent with the basic principles of QED [30]) parity-violating light-matter theories, we propose to focus rather on simpler parity-conserving couplings in combination with highly degenerate bare matter states. We expect that such a theoretical description is much more sensitive to chiral cavity-effects, since it does not rely on beyond dipolar couplings or strong external fields. Instead, degenerate states can be re-arranged (as in the case of the excited-state manifold of our quantum-ring example) and chiral symmetries may become favorable. Our proposal is further backed by recent theoretical results for linearly polarized cavities under collective vibrational strong coupling.[28] Those suggest the formation of a frustrated and highly-degenerate electronic ground state, which is prone to spontaneous symmetry breaking that impact on local molecular properties. [28] Accordingly, it seems plausible that a chiral cavity may also induce chiral symmetry breaking effects of chemical relevance. This will be the focus of future research efforts.

### v2

@@ IV Summary and conclusion
Nevertheless, our here-presented results suggest a remedy for this fundamental theoretical issue of chiral polaritonics. Instead of using numerically challenging and mathematically problematic (not every choice is consistent with the basic principles of QED [30]) parity-violating light-matter theories, we propose to focus rather on simpler parity-conserving couplings in combination with highly degenerate bare matter states. We expect that such a theoretical description is much more sensitive to chiral cavity-effects, since it does not rely on beyond dipolar couplings or strong external fields. Instead, degenerate states can be re-arranged (as in the case of the excited-state manifold of our quantum-ring example) and chiral symmetries may become favorable. Our proposal is further backed by recent theoretical results for linearly polarized cavities under collective vibrational strong coupling.[28, 36] Those suggest the formation of a frustrated and highly-degenerate electronic ground state, which is prone to spontaneous symmetry breaking that impact on local molecular properties. [28, 36] Accordingly, it seems plausible that a chiral cavity may also induce chiral symmetry breaking effects of chemical relevance. This will be the focus of future research efforts.
Supplementary Material
Additional simulation data is provided for the chiral cavity with an externally applied magnetic field.
Acknowledgments
This work was made possible through the support of the RouTe Project (13N14839), financed by the Federal Ministry of Education and Research (Bundesministerium fur Bildung und Forschung (BMBF)) and supported by the European Re- search Council (ERC-2015-AdG694097), the Cluster of Excellence “CUI: Advanced Imaging of Matter” of the Deutsche Forschungsgemeinschaft (DFG), EXC 2056, project ID 390715994 and the Grupos Consolidados (IT1249-19). The Flatiron Institute is a division of the Simons Foundation. C. M. Bustamante thanks the Alexander von Humboldt-Stiftung for the financial support from Humboldt Research Fellowship.